OpenStudy (lgbasallote):

Lgbariddle level 2: 2+5 = 12 3 + 8 = 30 4 + 6 = 36 5 + 8 = 60 6 + 4 = 54 7 + 6 = 84 8 + 9 =?

OpenStudy (lgbasallote):

Hint: Math related just as @Mimi_x3 wished

OpenStudy (anonymous):

121?

OpenStudy (lgbasallote):

nope try again :DDD

OpenStudy (anonymous):

128?

OpenStudy (lgbasallote):

lol are you guessing :P

OpenStudy (anonymous):

Nope my calculator was stupid

OpenStudy (anonymous):

2+5 = 12 => 2x5 + 2x1 =12 3 + 8 = 30 => 3x8 + 3x2 = 30 4 + 6 = 36 => 4x6 + 4x3 = 36 5 + 8 = 60 => 5x8 + 5x4 = 60 6 + 4 = 54 => 6x4 + 6x5 = 54 7 + 6 = 84 => 7x6 + 7x6 = 84 8 + 9 =? => 8x9 + 8x7 = 128

OpenStudy (anonymous):

Is it?

OpenStudy (lgbasallote):

ahh ahh nice..that was fast =))

OpenStudy (anonymous):

Cool~

OpenStudy (apoorvk):

When did you post this???? :-((((((((((

OpenStudy (anonymous):

38 minutes ago, as it said